OpenStudy (anonymous):

Multiply the fractions to get a simpler fraction! (1/2)·(3/1) A. 1/6 B. 2/3 C. 3/2 D. 4/3

OpenStudy (anonymous):

A

OpenStudy (jdoe0001):

\(\bf \cfrac{a}{b}\cdot \cfrac{c}{d}\implies \cfrac{a\cdot c}{b\cdot d}\)

OpenStudy (anonymous):

same thing

OpenStudy (the_fizicx99):